# APT-CGLP: Advanced Persistent Threat Hunting via Contrastive Graph-Language Pre-Training

### v1

@@ Abstract.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Security and privacy Intrusion detection systems
@@ 4.2.1. Encoders
where $\mathbf{h}_{n_{j}}\in\mathbb{R}^{d}$ is the $j$-th node feature, $|V_{i}|$ is the number of nodes in $G_{i}$, and $\theta_{G}$ denotes the parameters of GIN.
Text Encoder. We use BERT (bert) as the text encoder $f_{T}$, due to its effectiveness in capturing bidirectional causal dependencies among system entities within CTI reports. Each CTI report $T_{i}$ is first tokenized, and then passed through self-attention layers to produce token-level contextual representations:
@@
Multimodal Encoder. To learn semantic associations between provenance graphs and CTI reports, we introduce a multimodal encoder $f_{M}$, built on a transformer backbone. The encoder first applies self-attention over the token-level features $\mathbf{H}_{T_{i}}$ to capture intra-text dependencies, and subsequently utilizes cross-attention mechanisms to integrate structured node features $\mathbf{H}_{G_{i}}$. Such a design enables the model to learn detailed token-node interactions and achieve precise semantic alignment across modalities. The final multimodal representation is computed as:
(3) $\mathbf{z}_{m}=f_{M}(\mathbf{H}_{N_{i}},\mathbf{H}_{T_{i}};\theta_{M}),$
@@ 4.2.2. Training Strategy
Specifically, given a provenance graph-CTI report pair ($G_{i},T_{j}$), the unimodal encoders $f_{G}$ and $f_{T}$ are used to extract their respective overall representations. The graph embedding is derived by averaging the node embeddings: $\mathbf{z}_{G_{i}}=\frac{1}{|V_{i}|}\sum_{n\in V_{i}}\mathbf{h}_{n}$, while the CTI embedding $\mathbf{z}_{T_{j}}$ is extracted as the hidden state of its [CLS] token: $\mathbf{z}_{T_{j}}=\mathbf{h}_{t_{\text{[CLS]}}}$. We then compute the softmax-normalized similarity between the paired representations across a mini-batch as:
@@ 4.3. CTI Denoising Module
By guiding the LLM through this structured reasoning process, we obtain distilled intelligence that retains essential attack patterns while suppressing noise. Complete prompt templates and qualitative case studies are provided in Appendix C.
@@ 5. Evaluation
  • RQ2: What are the individual contributions of its core modules—training data augmentation, cross-modal training, and CTI denoising—to overall performance?
@@ 5.1. Experiment Settings
This section details datasets, evaluation protocol, and metrics. Implementation and hardware setup are provided in Appendix E.
@@ 5.1.2. Evaluation Protocol
  • CTI Corpus. The CTI corpus composes two sources: official DARPA reports and web-sourced reports from security vendors (e.g., MITRE (attck), VirusTotal (virustotal), Microsoft (microsoft)). To ensure quality and relevance, we apply regex-based filtering to an initial pool of 10,618 reports, retaining 5,183 CTI reports that contain detailed IoC and TTP descriptions. Keyword analysis confirms that these reports cover attack vectors such as phishing, living-off-the-land, and backdoor exploits. All CTI reports are processed through our denoising module to distill core insights, forming the external knowledge base for threat hunting.
Table 2 presents the ultimate testing dataset, comprising 2,402 subgraphs (1.2% malicious subgraphs) and 5,183 CTI reports.
@@ 5.3. RQ2: Ablation Studies
5.4. RQ3: Practicality of Retrieval Strategy
(a) Inference memory cost.
(b) Inference time cost.
Figure 5. Overhead of APT-CGLP for different $k$.
@@ 5.4. RQ3: Practicality of Retrieval Strategy
As reported in Table 5, APT-CGLP achieves AFRs over 90% across all four datasets, while maintaining consistently high TRRs. These results demonstrate its strong capability to eliminate spurious alerts without compromising coverage of genuine threats. Anomaly-based detectors (threatrace; kairos; aptkgl; magic) often exhibit elevated FPRs due to rigid thresholds and reliance on behavior baselines, exacerbating the problem of alert fatigue (nodoze)—a condition that delays or impairs incident response. APT-CGLP mitigates this issue by grounding alert filtering using external CTI knowledge.
@@ Acknowledgements
This work has been partially supported by the National Natural Science Foundation of China (Grant Nos. U22B2028, 62372410), the Zhejiang Provincial Natural Science Foundation of China (Grant No. LZ23F020011, LD22F020002), and the Zhejiang Province Leading Goose Program (Grant No. 2025C01013).
@@
Appendix A Activity Subgraph Sampling Algorithm
We developed a subgraph sampling algorithm to extract semantic-rich subgraphs from provenance graphs, serving as templates for generating corresponding CTI reports to build a cross-modal learning dataset. Additionally, we designed several guiding rules for sampling to ensure that the behavior semantics of the sampled graphs closely resemble real-world attack scenarios, thereby enhancing the learning effectiveness of APT-CGLP.
The algorithm employs a Depth-First Search (DFS) approach, as detailed in Algorithm 1. It takes as input a large provenance graph, sampling depth $k$, and a set of starting nodes. To make the subgraph closer to the attack scenarios, we first set $k=3$, enabling the DFS algorithm to capture activity paths with a maximum length of 7 (lines 3-17). Second, we limit the number of nodes in each sample to between 10 and 20 (line 18), reflecting the typical scale of entities involved in an attack activity. Finally, we restrict the subgraph to include all interactions types related to sockets, files, and processes (line 20), which are more representative of genuine attacks.
Input: Provenance graph pgpg; Sampling depth $k$; Starting nodes start_nodesstart\_nodes
@@ 1 $SG_{s}\leftarrow[]$
2
3foreach $v\in start\_nodes$ do
4  $visited\leftarrow Set(v)$
5  $stack\leftarrow[(v,0)]$
6  $sg\leftarrow\{\}$
7
8 while $len(stack)>0$ do
9     $(u,depth)\leftarrow stack.pop()$
10     $sg.add(u)$
11
12    if $depth<k$ then
13        $nei\_nodes\leftarrow pg.Neighbors(u)$
14        foreach $w\in nei\_nodes$ do
15           if $w\notin visited$ then
16              $visited.add(w)$
17              $stack.push(w,depth+1)$
@@ 6  $sg\leftarrow\{\}$
26 if $len(sg)>10$ & $len(sg)<20$ then
27     $edges\leftarrow sg.edges()$
28     if $edges\cap SOCKET\_EDGES\neq\emptyset$ & $edges\cap PROCESS\_EDGES\neq\emptyset$ & $edges\cap FILE\_EDGES\neq\emptyset$ then
29        $SG_{s}.add(sg)$
30
31
32  end if
33
34 end foreach
35
36return $SG_{s}$
@@ Appendix B CTI Generation Prompt
We harness LLaMA ’s in-context learning capabilities to construct synthetic training data, thereby enriching cross-modal training datasets of APT-CLIP. This approach facilitates the generation of desired CTI report formats by embedding specific demonstrations within the prompt. Notably, we customize 10 various in-context learning demonstrations to guide the LLM in generating sufficient diverse synthetic CTI reports, enabling the model to acquire robustness and generalization capabilities during training.
As illustrated in Figure 6, the prompt comprises three key components: task description, examples, and task input. The task description provides essential contextual information to enhance LLaMA ’s task-solving capabilities. The examples include structured triplet interactions derived from a provenance graph (as input), and corresponding activity descriptions in natural language form (as output), enabling LLaMA to internalize transformation rules. Finally, the task input represents the sequence of triplets that need to be processed. By leveraging the few-shot learning capability of LLaMA, we generate paired datasets of provenance graphs and their CTI reports in natural language forms, significantly enriching the dataset for cross-modal training.
@@ Appendix C CTI Denoising Prompt
We leverage LLaMA’s Chain-of-Thought techniques to systematically extract attack descriptions from unstructured CTI reports. This method is designed to reduce noise and generate coherent, actionable narratives for threat analysis. We present a case of CTI report denoising in Figure 8.
The prompt illustrated in Figure 7 comprises three main components: task description, multi-step processing instructions, and input data. The task description prompt provides LLaMA with crucial contextual information to improve comprehension and accuracy in processing CTI reports. The multi-step processing prompt splits the complex task into a series of simple sub-tasks. First, all key system entities within the CTI report, such as processes, files, and sockets, are identified. Then, we extract the interactions between these entities from the CTI report and categorize these interactions into limited categories (e.g., read, write, execute) to align with system audit frameworks. Finally, the extracted behaviors are organized into a coherent sequence, highlighting critical actions and events to maintain logical and chronological order. The output is a refined CTI report that accurately captures essential attack behaviors, providing concise and reliable narratives to support threat analysis.
@@ Appendix D Graph Isomorphism Network
We employ the Graph Isomorphism Network (GIN) (gin) as provenance graph encoder for its strong expressive power. GIN achieves the comparable discriminative capability as the Weisfeiler-Lehman graph isomorphism test in distinguishing different graph structures (behavior patterns), making it ideal for capturing complex semantics in provenance graphs.
The encoding process consists of aggregation, combination, and update phases, which recursively refine each node’s features by integrating semantic interactions from its neighbors. Initially, node features $h_{v}^{(0)}|v\in\mathcal{V}$ } are derived using the text encoder on their types and names, while edge features $e_{uv}|{u\in\mathcal{N}(v)}$} are initialized from specific interaction types. During the neighbor message aggregation phase, GIN aggregates features from all neighbors and their interactions to form aggregation messages:
$\text{agg}_{v}^{(k)}=\sum_{u\in\mathcal{N}(v)}g(h_{u}^{(k-1)},e_{uv})$
In the subsequent combination phase, layer-specific learnable parameter $\epsilon^{(k)}$ adjusts the weight of the node’s own features, integrating them with the aggregated messages as follows:
$\text{combined}_{v}^{(k)}=(1+\epsilon^{(k)})\cdot h_{v}^{(k-1)}+\text{agg}_{v}^{(k)}$
An MLP then applies nonlinear transformations to these combined features, resulting in updated node features:
$h_{v}^{(k)}=\text{MLP}^{(k)}(\text{combined}_{v}^{(k)})$
Following $k$ iterations of aggregation, a node’s representation effectively encapsulates the structural information of its $k$-hop network neighborhood. This iterative process ensures that each node’s features are enriched with contextual information from its surrounding subgraph. Upon updating node features across all layers, a global summation READOUT function aggregates these features into a comprehensive graph-level representation.
Figure 8. A case of CTI denoising, where subfigure (a) depicts a raw CTI report sourced from the web with line breaks removed, and subfigure (b) presents the CTI report after processing through the CTI denoising module.
Appendix E Environment and Implementation
All experiments were conducted on an Ubuntu 20.04 system with an Intel Xeon Gold 5128 CPU, two NVIDIA RTX 4090 GPUs, and 128 GB of RAM.
The implementation of APT-CLIP consists of approximately 8,000 lines of Python codes. We employed the first half of the BERT-Base model (bert-base) as the pre-trained text encoder and the latter half as the multimodal encoder, retaining all parameters at their default settings. The graph encoder is implemented as a 2-layer GIN model with the same dimensions as the text encoder. During training, the ratio of token masking was set as 15% in the MLM task, and one node was randomly masked in the MGM task. The AdamW optimizer was employed with a learning rate of $1\times 10^{-4}$ and a weight decay of 0.02. A cosine annealing scheduler was used to progressively reduce the learning rate to a minimum of $1\times 10^{-5}$ over 100 training epochs, with a warmup phase starting at $1\times 10^{-5}$ for the first 30 epochs. The weighting parameter $\alpha=0.7$ is set to balance the loss contributions during training. To mitigate overfitting, regularization and dropout strategies were implemented during model training. The vector database for storing and retrieving CTI report embeddings is based on FAISS (faiss). The optimal number of candidate CTI in the two-stage retrieval strategy was fixed at $k=10$, and the similarity threshold was set to $\lambda=0.6$. Furthermore, the open-source LLaMA3-8B model (llama3) was deployed locally to support synthetic training data generation and CTI denoising, owing to its low computational cost and improved security for handling sensitive data.

### v2

@@ Abstract.
copyright: ccjournalyear: 2026doi: 10.1145/3770854.3780275conference: Proceedings of the 32nd ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.1; August 09–13, 2026; Jeju Island, Republic of Koreabooktitle: Proceedings of the 32nd ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.1 (KDD ’26), August 09–13, 2026, Jeju Island, Republic of Koreaisbn: 979-8-4007-2258-5/2026/08ccs: Security and privacy Intrusion detection systems
@@ 4.2.1. Encoders
where $\mathbf{h}_{n_{j}}\in\mathbb{R}^{d}$ is the $j$-th node embedding, $|V_{i}|$ is the number of nodes in $G_{i}$, and $\theta_{G}$ denotes the parameters of GIN.
Text Encoder. We use BERT (bert) as the text encoder $f_{T}$, due to its effectiveness in capturing bidirectional causal dependencies among system entities within CTI reports. Each CTI report $T_{i}$ is first tokenized, and then passed through self-attention layers to produce token-level contextual embeddings:
@@
Multimodal Encoder. To learn semantic associations between provenance graphs and CTI reports, we introduce a multimodal encoder $f_{M}$, built on a transformer backbone. The encoder first applies self-attention over $\mathbf{H}_{T_{i}}$ to capture intra-text dependencies, and subsequently utilizes cross-attention mechanisms to integrate structured node semantics $\mathbf{H}_{G_{i}}$. Such a design enables the model to learn detailed token-node interactions and achieve precise semantic alignment across modalities. The final multimodal representation is computed as:
(3) $\mathbf{z}_{m}=f_{M}(\mathbf{H}_{G_{i}},\mathbf{H}_{T_{i}};\theta_{M}),$
@@ 4.2.2. Training Strategy
Specifically, given a provenance graph-CTI report pair ($G_{i},T_{j}$), the unimodal encoders $f_{G}$ and $f_{T}$ are used to extract their respective overall representations. The graph representation is derived via an attention pooling mechanism over node embeddings: $\mathbf{z}_{G_{i}}=\sum_{n\in V_{i}}\alpha_{n}\mathbf{h}_{n}$, where $\alpha_{n}$ is the attention weight of node $n$. The CTI representation $\mathbf{z}_{T_{j}}$ is extracted as the hidden state of its [CLS] token: $\mathbf{z}_{T_{j}}=\mathbf{h}_{t_{\text{[CLS]}}}$. We then compute the softmax-normalized similarity between the paired representations across a mini-batch as:
@@ 4.3. CTI Denoising Module
By guiding the LLM through this structured reasoning process, we obtain distilled intelligence that retains essential attack patterns while suppressing noise. Complete prompt templates are provided in Appendix C.
@@ 5. Evaluation
  • RQ2: What are the individual contributions of its core modules, namely training data augmentation, cross-modal training, and CTI denoising to overall performance?
@@ 5.1. Experiment Settings
This section describes the datasets, evaluation protocol, metrics, and the implementation and hardware configuration used in our experiments.
@@ 5.1.2. Evaluation Protocol
  • CTI Corpus. The CTI corpus composes two sources: official DARPA reports and web-sourced reports from security vendors (e.g., MITRE (attck), VirusTotal (virustotal), Microsoft (microsoft)). To ensure quality and relevance, we apply regex-based filtering to an initial pool of 10,618 reports, retaining 5,172 CTI reports that contain detailed IoC and TTP descriptions. Keyword analysis confirms that these reports cover attack vectors such as phishing, living-off-the-land, and backdoor exploits. All CTI reports are processed through our denoising module to distill core insights, forming the external knowledge base for threat hunting.
Table 2 presents the final testing dataset, which consists of 2,402 subgraphs and 5,183 CTI reports. Among them, 2,094 subgraphs are generated by Magic and are used for alert validation evaluation, while the remaining subgraphs (with 1.2% malicious) are generated by MEGR-APT and used for threat hunting evaluation.
@@
5.1.4. Environment and Implementation
All experiments were conducted on an Ubuntu 20.04 server equipped with an Intel Xeon Gold 5128 CPU, 128 GB of RAM, and two NVIDIA RTX 4090 GPUs.
The implementation of APT-CLIP comprises $\sim$5,000 lines of Python codes. We adopt the BERT-Base model (bert-base) as a pre-trained text encoder and fine-tune the last two layers. The graph encoder is implemented as a three-layer GIN network with its dimensionality aligned to that of the text encoder. A custom transformer-like model is designed as the multimodal encoder for cross-modal integration. During training, the token masking ratio for the MLM objective is set to 15%, and one node was randomly masked for the MGM objective. Optimization is performed using AdamW with a learning rate of $2\times 10^{-4}$ and a weight decay of 0.01. A cosine annealing schedule is applied over 100 epochs, decaying the learning rate to a minimum of $1\times 10^{-5}$. During the first 7 epochs, a warmup phase is employed. The loss-balancing weight is set to $\alpha=0.7$. To mitigate overfitting, regularization and dropout strategies were implemented during model training. The vector database for storing and retrieving CTI report embeddings is based on FAISS (faiss). The optimal number of candidate CTI in the two-stage retrieval strategy was fixed at $k=10$, and the similarity threshold was set to $\lambda=0.5$. Finally, the open-source LLaMA3-8B model (llama3) is deployed locally for synthetic training data generation and CTI report denoising, providing lower computational overhead and improved security when processing sensitive intelligence data.
@@ 5.4. RQ3: Practicality of Retrieval Strategy
As reported in Table 5, APT-CGLP achieves AFRs over 90% across all four datasets, while maintaining consistently high TRRs. These results demonstrate its strong capability to eliminate spurious alerts without compromising coverage of genuine threats. Anomaly-based detectors (threatrace; kairos; aptkgl; magic) often exhibit elevated FPRs due to rigid thresholds and reliance on behavior baselines, exacerbating the problem of alert fatigue (nodoze) that delays or impairs incident response. As benign behavior evolves over time, these systems often misclassify normal deviations as threats (depimpact). We mitigate this issue by grounding alert filtering using external CTI knowledge.
(a) Inference memory cost.
(b) Inference time cost.
Figure 5. Overhead of APT-CGLP for different $k$.
@@
6. Discussion
LLM Generalization: We tested various LLMs (e.g., Qwen2.5) and found that the proposed framework is not tied to a specific LLM. This is because the tasks performed in our framework (e.g., behavior extraction) are general, non-specialized tasks (bhusal2024secure; fieblinger2024actionable) that are well within the capabilities of most LLMs, while larger models typically perform better.
Hallucination Risk. The hallucinations risk of LLMs usage in our framework is partially limited by the closed, understanding-oriented nature of our tasks and the restricted graph scales during data synthesis. Future work will investigate this risk and potential mitigation strategies, such as consistency checking (chen2024inside).
Limitations. APT hunting evaluation requires datasets with paired audit data and attack reports, which, to our knowled ge, only DARPA datasets can satisfy. Other datasets that either lack complete audit records (e.g., StreamSpot (manzoor2016fast)) or do not provide detailed attack reports (e.g., LANL (turcotte2019unified)) are not suitable for this evaluation. Although our evaluation covers more than ten distinct attack types across different platforms, the scope of benchmarks remains limited. We leave the construction of more comprehensive and diverse APT datasets, encompassing broader attack behaviors and operating environments, as an important direction for future work.
@@ Acknowledgements
This work has been partially supported by the National Natural Science Foundation of China (Grant Nos. U22B2028, 62372410), the Key Research Program of Hangzhou (Grant No. 2025SZD1A56), the Zhejiang Provincial Natural Science Foundation of China (Grant No. LD22F020002), the Key Research Program of Shaoxing (Grant No. 2025B11004), the Zhejiang Provincial Natural Science Foundation of China (Grant No. LZ23F020011), and the Zhejiang Province Leading Goose Program (Grant No. 2025C01013).
@@
Appendix A Activity Subgraph Sampling
We design a subgraph sampling algorithm to extract semantic-rich subgraphs from provenance graphs, serving as templates for generating paired CTI reports to build a cross-modal learning dataset. As shown in Algorithm 1, it adopts a three-layer Breadth-First Search (BFS) strategy and takes as input a provenance graph $pg$ and a size range $[minN,maxN]$, and outputs a set of subgraphs $SG_{s}$.
To better approximate real attack scenarios, we design a progressive sampling process. In Layer 1 (lines 6-7), the algorithm samples process-type neighbors from socket nodes to capture initial attack entry points. In Layer 2 (lines 8-10), it performs balanced sampling across different node types through BFS expansion, ensuring structural diversity in the subgraph. In Layer 3 (lines 11-14), the subgraph is conditionally filled when its size falls below the minimum threshold, with additional nodes randomly sampled to reach a target size within the specified range. Finally, each sampled subgraph is validated to ensure it contains all node types and falls within the predefined size range (lines 15-16), guaranteeing representativeness of genuine attack patterns.
Input: Provenance graph pgpg; Size range $[minN,maxN]$
@@ 1 $SG_{s}\leftarrow[]$
2 $nei\_map\leftarrow PrecomputeNeighbors(pg)$
[…]
4
5foreach $v\in seed\_nodes$ do
6  $sg\leftarrow\{v\}$
  // Layer 1: Sample process neighbors
[…]
8  $sg.update(N_{1})$
9
 // Layer 2: Balanced sampling across node types
10  $N_{2}\leftarrow BFS(N_{1},nei\_map,sg)$
11  $N_{2}^{sampled}\leftarrow TypeBalancedSample(N_{2})$
12  $sg.update(N_{2}^{sampled})$
13
 // Layer 3: Conditionally fill subgraph
14  if $|sg|<minN$ then
15     $N_{3}\leftarrow BFS(N_{2}^{sampled},nei\_map,sg)$
16     $N_{3}^{sampled}\leftarrow RandomSample(N_{3},minN-|sg|,maxN-|sg|)$
17     $sg.update(N_{3}^{sampled})$
@@ 15     $N_{3}\leftarrow BFS(N_{2}^{sampled},nei\_map,sg)$
 // Validate subgraph
20  if $minN\leq|sg|\leq maxN$ & $CheckNodeTypes(sg)$ then
21     $SG_{s}.add(sg)$
22
23
@@
26return $SG_{s}$
@@ Appendix B CTI Generation Prompt
We harness LLaMA ’s in-context learning capabilities to construct synthetic training data, thereby enriching cross-modal training datasets. This approach facilitates the generation of desired CTI-like reports by offering specific demonstrations within the prompt. We customize 5 various in-context learning demonstrations to guide the LLM in generating sufficient diverse synthetic reports, enabling the model to acquire robustness and generalization capabilities during training. A prompt example is illustrated in Figure 6.
@@ Appendix C CTI Denoising Prompt
We employ LLaMA with Chain-of-Thought prompting to extract attack-relevant information from unstructured CTI reports. This method reduces noise and produces coherent, actionable narratives suitable for downstream threat analysis. A prompt example illustrated in Figure 7 comprises three main parts: task description, multi-step processing instructions, and input data.
@@ Appendix D Graph Isomorphism Network
We employ the Graph Isomorphism Network (GIN) (gin) as graph encoder for its strong expressive power. The encoding process comprises aggregation, combination, and update phases, which iteratively refine node features by integrating semantic signals from their neighbors. We initialize node features $h_{v}^{(0)}|v\in\mathcal{V}$ } using a text encoder applied to node types and names, while edge features $e_{uv}|{u\in\mathcal{N}(v)}$} are derived from their interaction types. At layer $k$, each node $v$ aggregates information from its neighbors through a learnable message function $g(\cdot)$:
$\text{m}_{v}^{(k)}=\sum_{u\in\mathcal{N}(v)}g(h_{u}^{(k-1)},e_{uv}).$
A trainable coefficient $\epsilon^{(k)}$ then controls the relative contribution of the node’s previous representation and the aggregated neighborhood message:
$\text{s}_{v}^{(k)}=(1+\epsilon^{(k)})\cdot h_{v}^{(k-1)}+\text{agg}_{v}^{(k)}.$
The updated node representation is produced by applying a layer-specific MLP:
$h_{v}^{(k)}=\text{MLP}^{(k)}(\text{s}_{v}^{(k)}).$
After $k$ iterations, each node embedding encodes information from its $k$-hop neighborhood, effectively capturing local structural and semantic context.